A particle P carries a charge e and moves under the influence of static magnetic field B(r) which exerts the force F × ev × B

on P, where v is the velocity of P . Show that P travels with constant speed.


1
Expert's answer
2021-05-09T18:17:00-0400

"F_1=evB,"

"F_2=ma=\frac{mv^2}{r},"

"F_1=F_2,"

"\frac{mv^2}{r}=evB,"

"v=\frac{eBr}{m}=const."